and 10 0.5 $1,061 calculations. Round your answers to 3 decima. owes
Answer #1

(a) Yield to Maturity = [C + (FV - PV)/n] / [(FV + PV)/2]

where, C = coupon payment = 8.6% of 1000/2 = $43

FV = Par Value = $1000

PV = Present Value of bond = $915

n = number of periods to maturity = 10*2 = 20 semiannual periods

=> YTM = [43 + (1000 - 915)/20] / [(1000 + 915)/2] * 2 = 0.0493* 2 or 9.86%

(b) Yield to Maturity = [C + (FV - PV)/n] / [(FV + PV)/2]

where, C = coupon payment = 5.7% of 1000/4 = $14.25

FV = Par Value = $1000

PV = Present Value of bond = $911

n = number of periods to maturity = 10*4 = 40 semiannual periods

=> YTM = [14.25 + (1000 - 911)/40] / [(1000 + 911)/2] *4= 0.0172*4 or 6.88%

(c) Yield to Maturity = [C + (FV - PV)/n] / [(FV + PV)/2]

where, C = coupon payment = 7.7% of 1000 = $77

FV = Par Value = $1000

PV = Present Value of bond = $1061

n = number of periods to maturity = 8

=> YTM = [77 + (1000 - 1061)/8] / [(1000 + 1061)/2] = 0.0673 or 6.73%